kekeman:

Describe how to determine the average rate of change between x = 1 and x = 3 for the function f(x) = 3x^3 + 1. Include the average rate of change in your answer.

mhanifa:

You just need to work out the values of f(1) and f(3), then divide the difference in f(3) and f(1) by the difference of 3 and 1

mhanifa:

That is:\[(f(3) - f(1))/(3 - 1)\]

kekeman:

Hmm ok thanks

mhanifa:

Did it work? Any help?

kekeman:

Yes would you like to see what i came up with?

mhanifa:

Yes, please

kekeman:

Here: https://snipboard.io/UZgydh.jpg

mhanifa:

Is this the one in your question?

mhanifa:

Can you find f(3) and f(1)?

mhanifa:

Substitute x with 3 or 1 in the given equation. f(3) = 3*3^3 + 1 f(1) = 3*1^3 + 1

kekeman:

I don't think so

mhanifa:

What do you mean?

kekeman:

So what i did was not correct?

mhanifa:

If you mean the attachment, it has nothing to do with your question

kekeman:

Ohhh i am sorry i sent you the wrong thing

kekeman:

https://snipboard.io/bVQZ3o.jpg

mhanifa:

This is the right one

kekeman:

Ok thank you so much

mhanifa:

No problem